OpenStudy (17bohtal):

Suppose that y varies inversely with x. Write an equation for the inverse variation. Y=4 when X=7 (a) x= y/3 (b) y= x/28 (c) y=3x (d) y= 28/x

OpenStudy (anonymous):

y = k/x 4 = k/7 k = 28 y = 28/x x = 28/y